OpenStudy (kamille):

hey,I need to solve the equation \[2\frac{ sinx }{ 2 }*\frac{ cosx }{ 2}*cosx=\frac{ 1 }{ 2 }\] \[cosxsinx*\frac{ cosx }{ 2 }=\frac{ 1 }{ 2 }|*2\] \[2sinxcosx*cosx=1\] \[\sin(2x)cosx=1\] Did I do anything wrongly? If no,what to do next?

OpenStudy (kamille):

@hartnn

OpenStudy (freckles):

What was the initial problem?

OpenStudy (kamille):

I need to solve the equation 2*sinx/2*cosx/2*cosx=1

OpenStudy (kamille):

sorry

OpenStudy (kamille):

2*sinx/2*cosx/2*cosx=1/2

OpenStudy (freckles):

\[2\sin(\frac{x}{2})\cos(\frac{x}{2})\cos(x)=\frac{1}{2} \\ \text{ or } \\ 2\frac{\sin(x)}{2}\frac{\cos(x)}{2}\cos(x)=\frac{1}{2}\]

OpenStudy (freckles):

\[2\sin(\frac{x}{2})\cos(\frac{x}{2})\cos(x)=\frac{1}{2} \\ \text{ Recall } \sin(x)=2 \sin(\frac{x}{2}) \cos(\frac{x}{2}) \\ \sin(x)\cos(x)=\frac{1}{2} \\ 2\sin(x)\cos(x)=1 \\ \sin(2x)=1 \]

OpenStudy (freckles):

That should help... The other thing I could probably give you a method that will lead to some extraneous answers.

OpenStudy (freckles):

sin(2x)=1 let u=2x find when sin(u)=1 (use unit circle)

OpenStudy (kamille):

oh,okay,thanks. I probably understood the problem incorrectly:)